Find the critical value

z Subscript alpha divided by 2zα/2

that corresponds to the confidence level

99​%.

z Subscript alpha divided by 2zα/2equals=

to two decimal places as​ needed.)
Answer #1

SOLUTION:

At 99% confidence level the z is ,

  = 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

Z/2 = Z0.005 = 2.58 ( Using z table )